OpenStudy (anonymous):

An equation of the line perpendicular to the graph of y = -2x + 4 that passes through the point at (-8,3)

OpenStudy (anonymous):

\[y - 3 = \frac{1}{2}(x + 8)\]\[y = \frac{x}{2} + 7\]